OpenStudy (anonymous):

a country's population in 1990 was 1554 million. in 2001 it was 159 million. estimate the population in 2005 using the exponential growth formula

OpenStudy (cwrw238):

P = P0 e^kt P = 159,000 , P0 = 154,000 and t = 2001 - 1990 = 11 plug these values into the formula and find the value of k

OpenStudy (cwrw238):

then you can use the formula to do the estimate